OpenStudy (anonymous):

What are the possible blood types among the children of parents that are IAIB and ii?

jagr2713 (jagr2713):

\[\huge{\color{purple}{\textbf{W}} \color{orange}{\cal{E}} \color{green}{\mathbb{L}} \color{blue}{\mathsf{C}} \color{maroon}{\rm{O}} \color{red}{\tt{M}} \color{gold}{\tt{E}} \space \color{orchid}{\mathbf{T}} \color{Navy}{\mathsf{O}} \space \color{OrangeRed}{\boldsymbol{O}} \color{Olive}{\mathbf{P}} \color{Lime}{\textbf{E}} \color{DarkOrchid}{\mathsf{N}} \color{Tan}{\mathtt{S}} \color{magenta}{\mathbb{T}} \color{goldenrod}{\mathsf{U}} \color{ForestGreen}{\textbf{D}} \color{Salmon}{\mathsf{Y}} \ddot \smile }\]

jagr2713 (jagr2713):

what types do you think @xBABYxGUBUx

OpenStudy (anonymous):

A, B, AB, O

jagr2713 (jagr2713):

ok um do you think AB would be one?

jagr2713 (jagr2713):

@Somy @Abhisar

OpenStudy (anonymous):

I think that it is all four

OpenStudy (abhisar):

For that, let's draw a punnet square first between \(\sf I^AI^B~and~ii\) 

OpenStudy (abhisar):

Now what blood groups do you think are formed?

OpenStudy (abhisar):

\(\sf I^Ai =? \\ I^Bi = ?\)